If the numerator of a fraction is doubled and the denominator is increased by 7, the value becomes 2/3; but if the denominator is doubled and the numerator is increased by 2, the value becomes 3/5. What is the original fraction?


1
Expert's answer
2022-04-07T05:40:59-0400

Let "\cfrac{x}{y}-" the original fraction, x - the numerator, y - the denominator.


"\begin{cases}
 \cfrac{2x}{y+7}=\cfrac{2}{3}\\
 \cfrac{x+2}{2y}=\cfrac{3}{5}
 \end{cases}"


"\begin{cases}
 6x=2y+14\\
5x+10=6y
 \end{cases}"


"\begin{cases}
 y=3x-7\\
5x-6(3x-7)+10=0
 \end{cases}"


"5x-18x+42+10=0\\
-13x+52=0\\
x=4\\
y=3\cdot4-7=5"


"\cfrac{4}{5}-" the original fraction.